If f(x)=(1/3)x + 3, then f-1(x) is?
I got 3x-3, but I got it wrong.

Found 2 solutions by nerdybill, palanisamy:
Answer by nerdybill(7384)

.
To find the inverse:
f(x)=(1/3)x + 3
can be rewritten as
y=(1/3)x + 3
Now, flip all the y's for x's and x's for y's:
x=(1/3)y + 3
Now, solve for y:
x-3 = (1/3)y
3(x-3) = y
3x -9 = y
.
Therefore,
f^-1(x) = 3x -9

Answer by palanisamy(496)

Let f(x)=(1/3)x + 3=y
x/3 = y-3
x = 3(y-3)
So, f-1(x) = 3(x-3)